OpenStudy (anonymous):

Impossible? Part 1 Is the following equation true? Part 2 Use complete sentences to explain the properties used in making your decision. log 3 x + 1/2 log 3 y – 2 log 3 z = log 3 x^4√y/z^2

OpenStudy (anonymous):

This is what I found but I don't understand it This equation is true. 4log[3] x + (1/2)log[3] y - 2log[3] z = log[3] x^4 + log[3] y^(1/2) - log[3] z^2 = log[3] (x^4)(sqrt(y)) - log[3] z^2 = log[3] [(x^4)(sqrt(y)) / z^2] Properties used: a log b= log b^a log a + log b = log (ab) log a - log b = log (a/b) a^(1/2) = sqrt(a)

OpenStudy (mertsj):

\[\log_{3}x+\frac{1}{2}\log_{3}y-2\log_{3}z=\log_{3}x^4\frac{y}{z^2} \]

OpenStudy (mertsj):

Is this the problem?

OpenStudy (anonymous):

x^4 should be on top with y

OpenStudy (mertsj):

\[\frac{1}{2}\log_{3}y=\log_{3}\sqrt{y} \]

OpenStudy (mertsj):

\[2\log_{3}z=\log_{3}z^2 \]

OpenStudy (mertsj):

So now we have: \[\log_{3}x=\log_{3}\sqrt{y}-\log_{3}z^2=\log_{3}\frac{x^4y}{z^2} \]

OpenStudy (mertsj):

Whoops...make that a plus sign

OpenStudy (mertsj):

log a + log b = log ab so we have: \[\log_{3}x \sqrt{y}-\log_{3}z^2=\log_{3}\frac{x^4y}{z^2} \]

OpenStudy (mertsj):

Now if we add \[\log_{3}z^2 \] to both sides we have: \[\log_{3}x \sqrt{y}=\log_{3}\frac{x^4y}{z^2}+\log_{3}z^2 \]

OpenStudy (mertsj):

But using log a + log b = log ab on the right side gives us: \[\log_{3}x \sqrt{y}=\log_{3}x^4y \]

OpenStudy (mertsj):

\[0=\log_{3}x^4y-\log_{3}x \sqrt{y} \]

OpenStudy (mertsj):

Doesn't look true to me.

OpenStudy (anonymous):

in the original problem did you make the x^4 so it was with square root/ z^2? @Mertsj

OpenStudy (anonymous):

*square root of y

OpenStudy (anonymous):